OpenStudy (anonymous):

Find g º f if f(x) = 2x - 1 and g(x) = x + 2.

OpenStudy (ddcamp):

g°f means g(f(x)) Replace every x in g(x) with f(x), then simplify.

OpenStudy (anonymous):

so itd be 2x-1=x+2

OpenStudy (ddcamp):

No. f(x) = 2x-1 g(x) = x + 2 g(f(x)) = f(x) + 2 = (2x-1) + 2

OpenStudy (anonymous):

ohh alright, thanks :D!!!

OpenStudy (anonymous):